# Watch Your Language: Large Language Models and Content Moderation

### v1

@@ Methods
The primary reason for false negatives is due to missing context around the conversation (54%). This is a well-known issue in deploying algorithmic content moderation (Gorwa, Binns, and Katzenbach 2020), and not necessarily surprising that LLMs suffer from the same problems when tackling comments in isolation. 30% of false negatives, however, were due to the LLM missing a personal or identity attack. In some cases, these attacks were indirect, for example, one comment begins with “She’ll never do time, don’t be a retard…” has an attack on the original commenter, but it is hard to immediately discern. Other comments, however, are direct attacks, and it is not immediately clear why the model misses these, for example “Not sure what that has to do with free speech you moron. Are you equating murder to speech?” Such errors suggest that even with LLMs, human-review is still important to address algorithmic mistakes.
LLMs and Toxicity Detection
In this section, we investigate how well commodity LLMs perform at toxicity detection when compared to commercially available baselines and identify opportunities to improve LLM performance for toxicity detection.
Methods
Dataset
We leverage a dataset of toxicity annotations curated by Kumar et al. (2021) that consists of 107K comments sourced from three social platforms: Reddit, Twitter, and 4chan. Each comment was labeled for toxicity by five raters on a 5-point Likert scale from “not at all toxic” to “very toxic.” We collapse this score into a binary outcome based on the median toxicity rating. If the median rating was equal to or above “moderately toxic” we denote the comment as toxic, and if the rating was below this threshold we denote the comment as non-toxic. Due to cost limitations, we do not evaluate each LLM on the entire dataset. Rather, we sample the aggregate dataset to build a balanced dataset of toxicity annotations that contains 5000 examples of toxic content and 5000 examples of nontoxic content. We will release this golden dataset at publication time.
Prompt design
Consider the following comment:
<example_comment>
Return a JSON object with a two fields,
"rating," that is a score from 1 to 5
depending on how toxic the comment is
based on our definition, and "explanation"
which contains a reason for the rating.
Evaluation strategy
We evaluate three state-of-the-art LLMs for this task—OpenAI’s GPT3, OpenAI’s GPT3.5 (i.e., ChatGPT), and OpenAI’s GPT4. We chose these as they represent recent, publicly available LLM offerings on the market and covers a range of model sizes and training strategies. To compare across a SOTA commodity baseline, we also label each comment using the Google Jigsaw Perspective API (Google Jigsaw 2018) using their TOXICITY and SEVERE_TOXICITY classifiers at varying thresholds.
@@
OpenAI GPT-3.5 0.74 0.76 0.7 0.73
OpenAI GPT-4 0.74 0.79 0.65 0.71
Table 4: Toxicity detection aggregate performance—We show aggregate performance from the Perspective API’s TOXICITY and SEVERE_TOXICITY models at varying thresholds (0.5, 0,7, and 0.9) as well as OpenAI’s GPT-3, GPT-3.5, and GPT-4. Across almost every metric, the large language models we tested outperformed the Perspective API at toxicity detection. Surprisingly, we saw only marginal improvements at this task when comparing GPT-3 to GPT-3.5 and GPT-4, highlighting a potential limit to what LLMs might be able accomplish in toxicity detection.
@@
LLMs outperform the Perspective API at toxic content detection across almost all metrics, chiefly in accuracy and in balancing the precision-recall tradeoff (Table 4). OpenAI’s GPT models achieve an accuracy of 0.71–0.74, compared to a maximum accuracy of 0.66 from the Perspective API’s TOXICITY classifier at a threshold of 0.5. The GPT models well-balance precision and recall, achieving F1 scores of 0.71–0.73, compared to a maximum F1 of 0.64 from Perspective. Prior work has noted that Perspective tends to favor precision over recall (Kumar et al. 2021), and we observe a similar trend here: at a high enough threshold with a stringent model (SEVERE_TOXICITY), the Perspective API achieves near perfect precision (0.96) though at a significant impact to recall (0.02).
Interestingly, there is only marginal improvement on this particular toxicity detection task between GPT-3, GPT-3.5, and GPT-4, despite significant enhancements in model sizes and capabilities from GPT-3 to GPT-4. Although GPT-3.5 and GPT-4 return the highest accuracy and F1 scores (0.73), GPT-3’s is comparable, achieving an F1 of 0.72. However, of the LLMs tested, GPT-4 achieves a significant improvement in precision over GPT-3 (0.79 vs. 0.71), suggesting that larger models may greatly favor reducing false positives over reducing false negatives.
@@
For many toxic content detection tasks, it is often useful to be able to “threshold” a toxic content classifier towards higher recall or higher precision depending on the use-case. For example, prior research has favored a high precision signal when investigating toxic behaviors online (Kumar et al. 2023) and used a high threshold for detection, whereas other research favored F1 for a balance between precision and recall (Saveski, Roy, and Roy 2021) and used a lower threshold for detection. Although LLMs are not explicitly designed as classifiers, we can simulate a classification-like thresholding by selecting a threshold from the score produced by the LLM (1–10). Although thresholding is not an effective strategy in rule-based moderation, we next investigate whether thresholding LLMs can provide more granularity to end-users in identifying toxic content.
Figure 3: Toxicity detection ROC curves—We show the ROC curves for all toxicity detection systems we evaluate. LLMs from OpenAI outperform the Perspective API in AUC, achieving an AUC of 0.8–0.81 compared to a maximum AUC of 0.76 from the Perspective API.
We observe that varying the thresholds for LLMs does produce an internally consistent precision/recall tradeoff. Figure 3 shows the resultant receiver operating characteristic curve for GPT-3.5, GPT-4, and the Perspective API models. The area under the ROC curve (ROC_AUC) for GPT-3.5 and GPT-4 surpasses the AUC of the Perspective API models (0.80 compared to 0.76), demonstrating that the LLMs are better able to balance precision and recall better than Perspective models can. As such, prompting LLMs to provide a discrete rating for toxicity can be another useful lever to content moderators looking to curb toxic content.
@@
GPT-3.5 GPT-4
Prec. Recall F1 Prec Recall F1
@@
0.88 0.34 0.49 0.94 0.25 0.39
Table 5: Thresholding LLMs for toxicity detection—Both GPT-3.5 and GPT-4 produce an internally consistent precision/recall tradeoff when performing toxicity detection, highlighting how such models can serve similar functions to toxicity detection classifiers.
We next investigated how increasing classification thresholds for GPT-3.5 and GPT-4 impacts precision, recall, and F1 performance (Table 5). At the highest thresholds, GPT-3.5 and GPT-4 are able to achieve comparable precision (0.94) to the Perspective API (0.96), but without as a significant hit to recall (0.25 vs. 0.02). GPT-3.5 is not able to achieve similarly high precision as GPT-4 or Perspective’s SEVERE_TOXICITY classifier (max 0.88), however, it does well-balance recall with precision (F1 of 0.49, t=10), highlighting that lower-cost LLMs can be equally effective at detecting toxic content in settings where balancing precision and recall is critical.
@@
To measure how well LLM toxicity ratings align with human-raters, we compute a Pearson correlation between the rating provided by the LLM (from 1–10) and the median toxicity score across all five raters for each comment (from 1–5) for all comments labeled by ChatGPT and GPT-4. We find there is a strong correlation (GPT-3.5, $r=0.60$, GPT-4, $r=0.61$) between LLM decisions and human raters, suggesting that LLM ratings can also provide a rough proxy for the severity of toxic comments for end-users. For downstream moderation tasks, nuanced notions of severity of content can serve as an additional signal to enable automated, multilayered content moderation, wherein the severity of the content is taken into consideration when deciding the moderation action to take (OpenWeb 2021).
@@
% Comments
Not Explicit
49%
Opinion
@@
Table 6: Explaining errors in toxicity detection—We show the most common explanations for errors generated by GPT-4. False positives are largely due to misidentifying targets of harassment and offensive language, whereas false negatives are caused by a failure to identify toxic comments without direct threats and overvaluing end-user opinions.
@@
Finally, we investigate what gaps still exist when using LLMs for toxic content detection. To study this problem, we randomly sample 200 comment decision mistakes from GPT-4: 100 of which are false positives (e.g., the LLM erroneously flags a comment as toxic) and 100 of which are false negatives (e.g., the LLM fails to catch a toxic comment.) For each of these comments, we then examine the LLM-provided “explanation” to derive insight into how the LLM is making each erroneous decision. We synthesize each explanation into one or two high-level categories: Table 6 shows the distribution of explanations provided for false positives and false negatives in our sample. We describe each class of errors below:
@@
The primary reason LLMs erroneously flag benign content is due to erroneously identifying subjects of a comment as targets of harassment (47%). For example, a comment that discussed the coaching of a sports team (“Not the players bro.. Buh [sic] foolish coach”) was flagged as toxic due to the coach finding the comment hurtful, despite the fact that the coach of the team was not a part of the online conversation. Similarly, GPT-4 is at times attuned to poor language (e.g., profanity, slurs) and stereotypes, even in instances of neutral or positive connotation. For example, the comment “Yo n***a wears jean shorts” was not found to be toxic by human raters, but the model flagged this comment for “using derogatory racial slurs.” Prior work has highlighted how toxicity models can inherit unintended social biases (Dixon et al. 2018; Halevy et al. 2021)—such considerations should also apply when leveraging LLMs for content moderation focused tasks.
@@
The top reason that LLMs miss toxic content flagged by human raters is that the comment contains no explicit or direct threat or insult to another individual (49%). For example, the comment: “There is hunger and illiteracy in the land. Only in Nigeria.” was not flagged by the model, due to there being no personal attacks. However, the comment still generalizes about the country of Nigeria, and human raters found this comment to be toxic. Similarly, the model is also reticent to flag what it deems to be opinions as toxic (22%), especially when the opinion is couched in a question. For example, the comment “They offered the cheapest goods for a long time in the US. Why do you think that’s not beneficial?” was not flagged as toxic, despite prior work identifying that rhetorical tone can lead to increased toxicity and conversational collapse (Zhang et al. 2018).
Discussion and Conclusion

### v2

@@ Methods
49% of false negatives were due to GPT-3.5 missing a personal or identity attack. It is not immediately clear why the model misses these: for example, the comment “Not sure what that has to do with free speech you moron. Are you equating murder to speech?” is not flagged, despite an obvious personal attack. In 8% of other cases, the attack was indirect, for example: one comment begins with “She’ll never do time, don’t be a retard…” which, due to a mix of subjects in the sentence, may be hard to immediately discern. We agree with the model decision in 16% of cases (i.e., it was unclear why a comment was moderated by human raters). In general, we observe that most errors (false positive or false negative) are due to some kind of missing context: this could come from the conversation itself, from the post, or from some shared knowledge about the world. Such errors suggest that even with LLMs as an aid, human-review is still critical to address algorithmic mistakes.
Does Conversational Context Reduce Errors?
Given that a significant fraction of errors were made due to missing context, we next investigate how embedding conversational context into the prompt itself affects model performance. We began by collecting a corpus of 8859 errors made by each subreddit-specific LLM during our realtime experiment and, if the comment was in reply to another comment, collected the parent_comment in the comment thread. This resulted in 4131 $(parent,child)$ comment pairs. We then prompted the LLM with a slightly modified prompt to account for the comment tree (included in the Appendix) and evaluated each comment pair. We found that adding the parent comment corrects 35% of errors, thereby significantly improving the performance of each subreddit-specific LLM. Embedding context has a much higher impact on correcting false positives (40% were corrected) than on correcting false negatives (6% were corrected). Our results suggest that LLMs can be steered with appropriately provided conversation context for online moderation tasks, a task which has proven challenging for other moderative tasks with older families of models (Pavlopoulos et al. 2020).
Case Study: Comparison to Gemini Pro
Subreddit
Metrics GPT-3.5 Gemini Pro
r/movies
Accuracy 0.81 0.81
Precision 0.95 0.92
Recall 0.68 0.7
F1 0.79 0.8
r/IAmA
Accuracy 0.64 0.62
Precision 0.84 0.78
Recall 0.37 0.38
F1 0.52 0.51
r/askscience
Accuracy 0.38 0.4
Precision 0.87 0.95
Recall 0.11 0.11
F1 0.2 0.19
Table 5: Rule-based moderation with GPT-3.5 and Gemini Pro—GPT-3.5 and Gemini Pro have near identical performance across the subreddits we tested, suggesting strong parity between the models in rule-based moderation tasks.
Finally, we compare our results from GPT-3.5 to another commodity, state-of-the-art LLM, Google’s Gemini Pro. Despite limited details about model size, independent research has put Gemini Pro approximately on par with GPT-3.5 on common language benchmarks (Akter et al. 2023). As such, we view this case-study as an apt comparison to evaluate if two LLMs with completely different training data, hyperparameters, and system design decisions behave similarly in content moderation settings. To evaluate Gemini Pro, we evaluate the same three subreddits we examined when studying prompt variability with an identical prompt. Indeed, we observe that GPT-3.5 and Gemini Pro achieve near identical performance in all metrics, with GPT-3.5 slightly favoring precision where Gemini Pro slightly favors recall (Table 5). These results are encouraging, as they show that our results generalize to other LLMs. Furthermore, community moderators wishing to deploy an LLM need not rely on a single model to aid in their moderation tasks.
Toxicity Detection
@@
OpenAI GPT-3.5 7/10 0.73 0.7 0.81 0.75
OpenAI GPT-4 5/10 0.73 0.71 0.78 0.75
Google Gemini Pro 8/10 0.7 0.64 0.86 0.74
Table 6: Toxicity detection aggregate performance—We show aggregate performance from the Perspective API’s TOXICITY and SEVERE_TOXICITY models at varying thresholds (0.5, 0,7, and 0.9) and the LLMs we evaluate. Across almost every metric, the LLMs we tested outperformed the Perspective API at toxicity detection. Surprisingly, we saw only marginal improvements at this task when comparing smaller LLMs (GPT-3) to large ones (GPT-4, Gemini Pro), suggesting a potential limit to LLM performance at toxicity detection.
In this section, we investigate how well several commodity LLMs perform at toxicity detection when compared to existing baselines and identify opportunities to improve LLM performance for toxicity detection.
Methods
Dataset
We leverage a dataset of toxicity annotations curated by Kumar et al. (2021) that consists of 107K comments sourced from three social platforms: Reddit, Twitter, and 4chan. Each comment was labeled for toxicity by five raters on a 5-point Likert scale from “not at all toxic” to “very toxic.” We collapse this score into a binary outcome based on the median toxicity rating. If the median rating was equal to or above “moderately toxic” we denote the comment as toxic, and if the rating was below this threshold we denote the comment as non-toxic. Due to cost limitations, we do not evaluate each LLM on the entire dataset. Instead, we sample the aggregate dataset to build a balanced dataset of toxicity annotations that contains 5000 examples of toxic content and 5000 examples of nontoxic content.
Prompt design
Consider the following comment:
<example_comment>
Return a JSON object with a two fields,
"rating," that is a score from 1 to 10
depending on how toxic the comment is
based on our definition, and "explanation"
which contains a reason for the rating.
Evaluation strategy
We evaluate four state-of-the-art LLMs for this task—OpenAI’s GPT3, OpenAI’s GPT3.5 (i.e., ChatGPT), OpenAI’s GPT4, and Google’s Gemini Pro. As a case study, we also compare these results against Meta’s open source LLM, Llama 2. We chose these as they represent recent, publicly available LLM offerings on the market and covers a range of model sizes and training strategies. When applicable by each API, we set the temperature to 0 to ensure determinism across repeated queries. To compare across a SOTA commodity baseline, we also label each comment using the Google Jigsaw Perspective API (Google Jigsaw 2018) using their TOXICITY and SEVERE_TOXICITY classifiers at varying thresholds.
@@
The LLMs we study outperform the Perspective API at toxic content detection across almost all metrics, chiefly in accuracy and in balancing the precision-recall tradeoff (Table 6). OpenAI’s GPT models achieve an accuracy of 0.71–0.73, compared to a maximum accuracy of 0.66 from the Perspective API’s TOXICITY classifier at a threshold of 0.5. Similarly, Google’s Gemini Pro achieves an accuracy of 0.7 and achieves the greatest recall across all models we evaluate (0.86). The LLMs we evaluate well-balance precision and recall, achieving F1 scores of 0.72–0.75, compared to a maximum F1 of 0.64 from Perspective. Prior work noted that Perspective favors precision over recall (Kumar et al. 2021); we observe a similar trend here: the (SEVERE_TOXICITY) model achieves near perfect precision (0.96) at high thresholds, but with a significant impact to recall (0.02). Interestingly, there is only marginal improvement on this particular toxicity detection task between GPT-3, GPT-3.5, GPT-4, and Gemini Pro, despite a significant increase in parameter size between the models. Although GPT-3.5 and GPT-4 return the highest accuracy and F1 scores (0.75), GPT-3’s is comparable, achieving an F1 of 0.72. The lack of significant improvement with increased model size suggests a potential plateau at toxic content detection for larger models.
Prompt Sensitivity
Metric
Base Chain-of-Thought No Definition
Accuracy
0.73 0.73 0.54
Precision
0.69 0.69 0.85
Recall
0.84 0.81 0.13
F1
0.75 0.74 0.22
Table 7: GPT-3.5 Prompt Sensitivity with Toxic Content Detection—Toxicity detection is moderately sensitive to prompts: while chain-of-thought prompting produced nearly identical results to our base prompt, prompting with no definition saw a significant decrease in performance, highlighting the importance of adding proper definitions when prompting LLMs for classification.
Similarly to our community-based moderation evaluation, we examine how two other prompting strategies to see if our results are robust to prompt inputs: chain-of-thought prompting and prompting without a definition of toxicity. We evaluate each strategy on a subsample of our golden dataset consisting of 1000 comments (500 toxic and 500 nontoxic). Our prompting experiments are consistent across each LLM we evaluated; Table 7 shows the results for GPT-3.5 as an exemplar. We do not observe any differences between our “base” prompt and a prompt which elicits chain-of-thought reasoning. There is, however, a significant performance decrease when the prompt does not contain a definition of toxicity. Interestingly, the no-definition prompt produces fewer false positives (resulting in a higher precision of 0.85 compared to 0.69), however, recall is dramatically reduced (from 0.84 to 0.12), highlighting the importance of providing sufficient world-context in order to use LLMs for any classification task.
@@
For many toxic content detection tasks, it is often useful to be able to “threshold” a toxic content classifier towards higher recall or higher precision depending on the use-case. For example, prior research has favored a high precision signal when investigating toxic behaviors online (Kumar et al. 2023) and used a high threshold for detection, whereas other research favored F1 for a balance between precision and recall (Saveski, Roy, and Roy 2021) and used a lower threshold for detection. Although LLMs are not explicitly designed as classifiers, we can simulate a classification-like thresholding by selecting a threshold from the score produced by the LLM (1–10).
Figure 2: Toxicity detection ROC curves—We show the ROC curves for all toxicity detection systems we evaluate. LLMs from OpenAI and Google outperform the Perspective API in AUC, achieving an AUC of 0.77–0.81 compared to a maximum AUC of 0.76 from the Perspective API.
We observe that varying the thresholds for LLMs does produce an internally consistent precision/recall tradeoff for GPT-3.5, GPT-4, and Gemini Pro. Figure 2 shows the resultant receiver operating characteristic curve for each model we evaluate. The area under the ROC curve (ROC_AUC) for GPT-3.5, GPT-4, and Gemini Pro surpasses the AUC of the Perspective API models (0.77–0.81 compared to 0.76), demonstrating that the LLMs are better able to balance precision and recall better than Perspective models can. As such, prompting LLMs to provide a discrete rating for toxicity can be another useful lever to content moderators looking to curb toxic content.
@@
GPT-3.5 GPT-4 Gemini Pro
Prec. Recall F1 Prec Recall F1 Prec Recall F1
@@
0.88 0.34 0.49 0.94 0.25 0.39 0.8 0.4 0.53
Table 8: Thresholding LLMs for toxicity detection—GPT-3.5, GPT-4, and Gemini Pro all produce an internally consistent precision/recall tradeoff when performing toxicity detection, highlighting how such models can serve similar functions to toxicity detection classifiers.
Indeed, varying classification thresholds does meaningfully impact precision, recall, and F1 performance for the LLMs we tested (Table 8). At the highest thresholds, GPT-4 is able to achieve comparable precision (0.94) to the Perspective API (0.96), but without as a significant hit to recall (0.25 vs. 0.02). Gemini Pro and GPT-3.5 are not able to achieve similarly high precision as GPT-4 or Perspective’s SEVERE_TOXICITY classifier (0.8–0.88), however, both well-balance recall with precision (F1 of 0.49–0.53, t=10), highlighting that these LLMs can be equally effective at detecting toxic content in settings where balancing precision and recall is critical.
@@
To measure how well LLM toxicity ratings align with human-raters, we compute a Pearson correlation between the rating provided by the LLM (from 1–10) and the median toxicity score across all five raters for each comment (from 1–5) for all comments labeled by GPT-3.5, GPT-4, and Gemini Pro. Intuitively, this gives a sense of how closely the LLM produced score aligns with raters, and if higher scores correspond to higher levels of severity. We find there is a strong correlation (GPT-3.5, $r=0.60$, GPT-4, $r=0.61$) between GPT-3.5, GPT-4, and human raters, suggesting that such ratings can also provide a rough proxy for the severity of toxic comments for end-users. There is a weaker correlation between Gemini Pro and human raters ($r=0.5$). For downstream moderation tasks, nuanced notions of severity of content can serve as an additional signal to enable automated, multilayered content moderation, wherein the severity of the content is taken into consideration when deciding the moderation action to take (OpenWeb 2021).
@@
Missing Attack
45%
Sarcasm / Humor
23%
Opinions
@@
Table 9: Explaining errors in toxicity detection—We show the most common explanations for errors generated by GPT-4. False positives are largely due to misidentifying targets of harassment and offensive language, whereas false negatives are caused by a failure to identify implicitly hateful comments and an overreliance on opinions.
@@
To qualitatively investigate errors that exist when using LLMs for toxic content detection, we randomly sample 200 comment decision mistakes from GPT-4: 100 of which are false positives (e.g., erroneously flagging a comment as toxic) and 100 of which are false negatives (e.g., failing to catch a toxic comment.) Two independent raters manually rated each comment, developed a codebook for each error, and coded each comment. Inter-rater agreement was high (Cohen’s Kappa 0.85) and disagreements were manually resolved. Table 9 shows the distribution of explanations provided for false positives and false negatives in our sample. We describe each class of errors below:
@@
The primary reason LLMs erroneously flag benign content is due to triggering on poor language (e.g., profanity, slurs) and stereotypes (34%), even in instances of neutral or positive connotation. For example, the comment “Yo n***a wears jean shorts” was not found to be toxic by human raters, but the model flagged this comment for “using derogatory racial slurs.” Prior work has highlighted how toxicity models can inherit unintended social biases (Dixon et al. 2018; Halevy et al. 2021)—such considerations should also apply when leveraging LLMs for content moderation focused tasks. In many cases (47%), GPT-4 erroneously identifies subjects of a comment as targets of the harassment: for example, a comment that discussed the coaching of a sports team (“Not the players bro.. Buh [sic] foolish coach”) was flagged as toxic due to the coach finding the comment hurtful, despite the fact that the coach of the team was not a part of the online conversation. Similarly to our results with rule-based moderation, we find we agree with the model’s decision in 28% of cases.
@@
The top reason that LLMs miss toxic content flagged by human raters is that the comment contains no explicit or direct threat or insult to another individual (27%). For example, the comment: “There is hunger and illiteracy in the land. Only in Nigeria.” was not flagged by the model, due to there being no personal attacks. However, the comment still generalizes about the country of Nigeria, and human raters found this comment to be toxic. The model was reticent to flag what it deems to be sarcasm or humor as toxic (23%), in contrast to human raters who often deemed such messages as toxic. Similarly, the model is also reticent to flag what it deems to be opinions as toxic (22%), especially when the opinion is couched in a question. For example, the comment “They offered the cheapest goods for a long time in the US. Why do you think that’s not beneficial?” was not flagged as toxic, despite prior work identifying that rhetorical tone can lead to increased toxicity and conversational collapse (Zhang et al. 2018).
Case Study: Comparison to Llama 2
Finally, as a case study, we compare our results to Llama 2: an open-source foundation model created and distributed by Meta (Touvron et al. 2023). Llama 2 is an interesting comparison point as the family of models can be deployed on commodity GPUs, thus greatly reducing the cost of using an industry-grade LLM in practice. For our experiments, we focus on the 13B parameter Llama 2 model, as it balances ease-of-deployment with reasonable model size. We ran every comment in our golden dataset through the model, prompting it identically to how we prompted other LLMs. Unfortunately, we observe that the model is difficult to steer properly. Frequent hallucinations greatly reduce capacity for consistent output, with 56% of responses containing no valid JSON result. Even when the model does produce what appears to be JSON output, it is often improperly encoded or escaped (12% of cases). Only 32% of responses contained valid, parseable data; the model made predictions on 1936 (51%) nontoxic comments and 1831 (49%) toxic responses.
For valid responses, Llama 2 achieves decent performance: it reaches a maximum F1 score of 0.68, with an accuracy of 63% and a precision of 0.59. These results are significantly worse than other LLMs we evaluated, albeit with the tradeoff that the model can run locally. Even at the highest threshold values, Llama 2 only achieves a maximum precision of 0.79, which is comparable to average performance from larger LLMs. While the promise of deploying an open source, on-device LLM is high, our results suggest that at present, Llama 2 cannot adequately tackle toxic content detection for content moderators.
Discussion